OpenStudy (imyoursquirtle):

Find the surface area of a cylinder with diameter 22 in and height 15 in.

OpenStudy (sidsiddhartha):

surface area of a cylinder is \[2*\pi*r ^{2}+2*\pi*r*h\] ur r=22/2=11 and h=15 now just put those values

OpenStudy (imyoursquirtle):

1796.99?

OpenStudy (sidsiddhartha):

teah thats correct